OpenStudy (anonymous):

Solve for \[c^2 + 5d^2\]: \[a^2 + 5b^2 = 10 \\ bc - ad = 5 \\ ac + 5bd = 15\]

OpenStudy (anonymous):

from second equation \( 5(bc-ad)^2=125 \) and from third u have \( (ac+5bd)^2=225 \) expand and then add this 2 equation and tell me what u get?

OpenStudy (anonymous):

\[5(bc)^2 + 5(ad^2) - 10abcd = 125 \\ (ac)^2 + 25(bd)^2 + 10abcd= 225\] Is this correct so far?

OpenStudy (anonymous):

Quite right ; add them

OpenStudy (anonymous):

\[0(ac)^2 + 5(bc)^2 + 5(ad^2) + 0(bd)^2 - 10abcd = 125 \\ (ac^2) + 0(bc)^2 + 0(ad)^2 + 25(bd)^2 + 10abcd = 225\]\[(ac^2) + 5(bc)^2 + 5(ad^2) + 25(bd)^2 = 350\] Is this right so far? In the beginning, how did you know to use these two equations and also to square and multiply them by a constant?

OpenStudy (anonymous):

well thats experience of problem solving; and u will get this skill through involvement such problems ; i did that because of form of a^2+5b^2=10

OpenStudy (anonymous):

c^2(.........

OpenStudy (anonymous):

I managed to expand it to\[a^2c^2 + 5b^2c^2 +5a^2d^2 +25b^2d^2 = 350\]I am rusty with my factoring...how do I take out both \[a^2+5b^2\] at the same time?

OpenStudy (anonymous):

Oh wait, let me try something else

OpenStudy (anonymous):

\[a^2(c^2 + 5d^2) + 5b^2(c^2 + 25d^2)\]\[(a^2 + 5b^2)(c^2 + 5d^2)(c^2 + 25d^2)\] Am I allowed to do this?

OpenStudy (anonymous):

be carefull thats \(a^2(c^2+5d^2)+5b^2(c^2+5d^2) \)

OpenStudy (anonymous):

Oh right, I forgot to factor out the '5' from '25'

OpenStudy (anonymous):

I see now...\[(a^2 + 5b^2)(c^2 + 5d^2)^2\]I can do this?

OpenStudy (anonymous):

I don't think the square belongs there, actually

OpenStudy (anonymous):

So \[(10)(c^2 + 5d^2) = 350\]and I need to find \[(c^2 + 5d^2)\] so divide, right?

OpenStudy (anonymous):

omg mukushla you are a genius...I wish I could know how to solve this without help (the squaring and multiplying in the beginning)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Answer: 35

OpenStudy (anonymous):

Quite right

OpenStudy (anonymous):

u will get there just keep on trying and solving many many problems like this ;)